OpenStudy (anonymous):

I need help studying for an oral exam , these are the topics she will be testing me on . Transformations Angle Measures Naming Polygons Symmetry Rotational Symmetry I need help on explaining rotational symmetry and angle measures . Thanks

OpenStudy (anonymous):

If interested, Come on g-mail.

OpenStudy (anonymous):

well symetry is easy it wether you can cut a figure evenly and all the peices look alike

OpenStudy (anonymous):

rotationl symmetry is a bit harder

OpenStudy (anonymous):

it is how many times the figure look the same after rotating it 45 degrees

OpenStudy (anonymous):

u r in which class/std?

OpenStudy (anonymous):

transformations is easy too

OpenStudy (anonymous):

there are translations, reflections, dilation, and rotations

OpenStudy (anonymous):

translations are simply moving the shape

OpenStudy (anonymous):

reflections are the flipped figure of the original over a line

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

dilations are changing the size, making it larger or smaller

OpenStudy (anonymous):

just wait.

OpenStudy (anonymous):

and rotations are rotated figures of the original

OpenStudy (anonymous):

all these have formulas for certain transformations

OpenStudy (anonymous):

angle measures, thats like sine cosine tangent, cosecant and sencant, and cotangents

OpenStudy (anonymous):

as well as the special triangles

OpenStudy (anonymous):

as well as the special triangles

OpenStudy (anonymous):

as for naming polygons, polygons are classifying by the number sides they have.

OpenStudy (anonymous):

for example pentagon hexagon heptagon octagon nonagon

OpenStudy (anonymous):

of course triangle and square before pentagon

OpenStudy (anonymous):

anything else in specific?

OpenStudy (anonymous):

No thats it that is what she will be asking me on .

OpenStudy (anonymous):

okay well good luck :)

OpenStudy (anonymous):

Translations- Up-Down f(x)=x2 • f’(x)= f(x)+C=x2+C [UP] • f’(x)= f(x)-C=x2-C [DOWN] Graph for x2 + 3 looks like this: To move a function up, you add outside the function: f(x)+C is f(x) moved up C units. Moving the function down works the same way; f(x)-C is f(x) moved down C units. Left-Right f(x)=x2 • f’(x)= f(x+C)=(x+C)2 [LEFT] • f’(x)= f(x-C)=(x-C)2 [RIGHT] Graph for (x + 3)2 looks like this: To shift a function left, add inside the function's argument: f(x + C) gives f(x) shifted C units to the left. Shifting to the right works the same way; f(x – C) is f(x) shifted C units to the right. Reflections:- Up-Down f(x)=x2 • f’(x)= -f(x)=-(x)2 [X-AXIS MIRROR REFLECTION] Graph for –x2 looks like this: -f(x) is the x-axis mirror image of f(x). Left-Right f(x)=x3 • f’(x)= f(-x)=(-x)3 [Y-AXIS MIRROR REFLECTION] Graph for –x3 looks like this: f(–x) is the y-axis mirror image of f(x). Functions with symmetry graph. Even Function- • f is even function if f(-x)=f(x) • Even functions are symmetry about y-axis. • Cosine is an even function. Odd Function- • f is odd function if f(-x)=-f(x) • Odd functions always pass through origin. • Odd functions graph remains unchanged after rotation of 1800 about the origin. • Sine is an odd function.